Solve the equation: 4x + 13 - 2 = 2 Answer: =
Answer #1

2 The+ 13 J 4c + 3 rt ? 4e4 13 =(E+ 922 4c+3 с”+ 4 +4 С.: 6+) а?+ь? + 2sb] + 4c+4 - Hс 13 +4 + 4 =( 3 съ= 13 - 24 2 2 чCs k2